OpenStudy (anonymous):

what's the name for 2 pi

OpenStudy (tkhunny):

"Steve"?

OpenStudy (anonymous):

NOO!!!!!!

OpenStudy (anonymous):

lol

OpenStudy (tkhunny):

Seriously, it's just \(2\pi\). Is there a need to call it something else?

OpenStudy (anonymous):

yes that's correct....look it up

OpenStudy (anonymous):

Tau

jimthompson5910 (jim_thompson5910):

from these pages, it looks like tau, but not 100% on that myself http://constitutionclub.org/2011/07/02/even-math-is-changing/ http://math-blog.com/2010/06/28/forget-pi-here-comes-tau/

OpenStudy (anonymous):

Also 'full circle' / '1 complete period of sine or cosine function,' etc.

OpenStudy (anonymous):

"360º" and so on, and so forth . . .

OpenStudy (tkhunny):

I still like "Steve".

OpenStudy (anonymous):

Steve is a good name too. I think Tau might have an advantage - still being a Greek letter and all that. You know how mathematical constants can be . . .

OpenStudy (tkhunny):

Why do we use \(2\pi r\) when \(\pi d\) is perfectly satisfactory - except that it is not a natural result of the calculus derivation. \(\dfrac{d}{dr}\pi r^{2} = 2\pi r\). It would be counter productive to replace either \(2r = d\) or \(2\pi = \tau\)

OpenStudy (anonymous):

How about angles around the unit circle? Quarter circle: π/2 or τ/4?

OpenStudy (anonymous):

or total radian of the circle (?)